OpenStudy (anonymous):

. In a market survey, 20% opted for product A whereas 60% opted for product B. The remaining individuals were not certain. If the difference between those who opted for product B and those who were uncertain was 720, how many individuals were covered in the survey ?

OpenStudy (anonymous):

let x be the number of people in the survey. 20% were uncertain so 0.60 x - 0.2x = 720 0.40x = 720 x = 720 / 0.4 = 1800